Please help me solve this problem: How many kilograms of 50 percent silver alloy and 10 percent silver alloy must be combined to obtain 24 kilograms of a 25 percent of silver alloy?
Answer by lwsshak3(11628)

How many kilograms of 50 percent silver alloy and 10 percent silver alloy must be combined to obtain 24 kilograms of a 25 percent of silver alloy?
***
let x=amt of 50% silver alloy to be combined
24-x=amy of 10% silver alloy to be combined
50%x+10%(24-x)=25%*24
.50x+2.4-.10x=6
.40x=3.6
x=9
24-9=15
amt of 50% silver alloy to be combined=9 kg
amy of 10% silver alloy to be combined=15 kg